# Event-Customized Image Generation

### v1

@@ 4.1 Experimental Setup
Evaluation Benchmarks. In order to provide sufficient and suitable conditions for both quantitative and qualitative comparisons on this new task, we collect two new benchmarks [^1] Due to the limited space, more details are left in the Appendix. . 1) For quantitative evaluation, we present SWiG-Event, a benchmark derived from the SWiG (Pratt et al., 2020) dataset, which comprises 5,000 samples with various events and entities, i.e., 50 kinds of different actions, poses, and interactions, where each kind of event has 100 reference images, and each reference image contains 1 to 4 entities with labeled bounding boxes and nouns. 2) For qualitative evaluation, we present Real-Event, which comprises 30 high-quality reference images from the HICO-DET (Chao et al., 2015) and the internet with a wide range of events and entities (e.g., animal, human, object, and their combinations). We further employ the Grounded-SAM (Kirillov et al., 2023; Ren et al., 2024) to extract the mask of each entity.
Baselines. To evaluate the effectiveness of our method, we compared it with several kinds of state-of-the-art baselines. For conditioned text-to-image generation baselines, we compared with the training-based method ControlNet (Zhang et al., 2023) and the training-free method BoxDiff (Xie et al., 2023). For localized editing baselines, we compared with training-free methods PnP (Tumanyan et al., 2023) and MAG-Edit (Mao et al., 2023). For customization baselines, we compared with training-based methods Dreambooth (Ruiz et al., 2023) and ReVersion (Huang et al., 2023).
Implementation Details. We use Stable Diffusion v2-1-base as the base model for all methods, and images are generated with a resolution of 512x512 on a NVIDIA A100 GPU1.
@@ 4.2 Quantitative Comparisons
In this subsection, we compare our method with conditional text-to-image generation baselines ControlNet (Zhang et al., 2023) and BoxDiff (Xie et al., 2023) on the SWiG-Event benchmark.
Retrieval-based Evaluation. Each reference image in SWiG-Event contains reference entities together with labeled event class, bounding boxes, nouns, and their corresponding masks. Specifically, we construct the target prompt as a list of reference entity nouns, i.e., we ask all the methods to reproduce the event of the reference image with the same reference event and same reference entities. Additionally, ControlNet takes the semantic map merged from the masks as the layout condition, and BoxDiff takes the bounding boxes with labeled entity nouns as the layout condition1. To evaluate the customization quality of 5,000 target images, we retrieved each target image for its corresponding reference image across all the 100 reference images that have the same reference event class. Specifically, we extracted the image feature of each image through a pre-trained CLIP (Radford et al., 2021) visual encoder and calculated the cosine similarities for image retrieval.
Results. As shown in Table 1, our FreeEvent achieves superior retrieval performance over both ControlNet and BoxDiff. Specifically, the Recall@1 of FreeEvent is significantly higher than that of the baselines, and it even surpasses the Recall@10 of the baselines. This demonstrates that the target images generated by FreeEvent better preserve the characteristics of the reference event and entity, showcasing a stronger capability for event customization.
Model Recall@1 Recall@5 Recall@10
ControlNet (Zhang et al., 2023) 10.64 26.12 36.82
BoxDiff (Xie et al., 2023) 8.60 22.48 32.08
Free-Event (ours) 41.12 63.02 72.74
Table 1: Performance of our model and state-of-art conditional text-to-image generation models on SWiG-Event. The Recall@k represents that among the top-k images with the highest similarity to the target image, its corresponding reference image is included.
@@ 4.3 Qualitative Comparisons
We compare FreeEvent with a wide range of state-of-the-art baselines on the Real-Event benchmark, including the conditioned text-to-image generation methods ControlNet (Zhang et al., 2023) and BoxDiff (Xie et al., 2023), the localized image editing methods PnP (Tumanyan et al., 2023) and MAG-Edit (Mao et al., 2023), the image customization methods Dreambooth (Ruiz et al., 2023) and ReVersion (Huang et al., 2023).
@@
Results. As shown in Figure 4, we can observe: 1) Conditional text-to-image generation models ControlNet and BoxDiff can only maintain the rough layout of each entity and struggle to capture the detailed action, pose, or interaction between different entities. And they both failed to match the generated entity with the desired target prompt. 2) For localized image editing methods PnP and MAG-Edit, while they can capture the reference event, they both struggle to accurately generate the target entities, and suffer from severe appearance leakage between each target entity (e.g., orange and strawberry in row four, tiger and lion in row six), and sometimes even failed to edit and output the original content. 3) The subject-customization model Dreambooth and the relation-customization model ReVersion both failed to generate satisfying results. As discussed before, these training-based methods require multiple reference images and are unable to learn the specific event when facing only one reference image. 4) Obviously, our FreeEvent successfully achieves the customization of various complex events with novel combinations of target entities. Meanwhile, the ControlNet and the localized image editing models tend to generate the target entities strictly matching the mask of their corresponding reference entities (e.g., bird in row three), which appears very incongruous. On the contrary, the entities generated by FreeEvent not only match the layout of the reference entity but also keep it harmonious. After all, while we use the reference entity mask to guide the generation of each target entity, the cross-attention guidance focuses on directing the overall layout of each target entity and does not restrict their detailed appearance, allowing for a more diverse generation of target entities1.
@@ 4.4 Ablations
Influence of Different Target Prompts. Noteablly, in our paper, the target prompt only contains the nouns of the target entities, we then run the ablations to analyze the influence of different descriptions (i.e., verb, background, style) in the target prompt for event customization.
@@
Figure 5: Ablations of the proposed paths and the target prompt. The “guidance" and “regulation" denote the cross-attention guidance and cross-attention regulation in the entity switching path, respectively. The “injection" denotes the event transferring path.
Combination of Event and Subject Customization. We further validate the ability of our framework to combine with subject customization methods to generate target entities with user-specified subjects, i.e., represented by identifier tokens. For an easy exploration, we took DreamBooth (Ruiz et al., 2023) to learn identifier tokens for each subject and replaced the Stable Diffusion models in Figure 2 with the fine-tuned one.
Results. As shown in Figure 6, FreeEvent can effectively generate various given subjects in specific events, showing great potential for Event-Subject customization.
@@ 4.5 User Study
Setting. We conducted user studies on Real-Event to further evaluate the effectiveness of FreeEvent. Specifically, we invited 10 experts and gave them a reference image, a target prompt, and seven target images generated by different models. They are asked to choose the three target images that they believe demonstrate the best results in event customization, taking into account the generation effects of the events and entities, as well as the overall coherence of the images. We prepared 30 trials and asked the experts to give their judgments. The target image which got more than six votes is regarded as human judgment.
@@
Model Ours ControlNet BoxDiff PnP MAG-Edit DreamBooth ReVersion
HJ 30 12 2 19 9 1 0
Table 2: Results of the user studies on the Real-Event. “HJ" denotes the count of human judgment.
@@ 5 Conclusion
In this paper, we proposed a new image generation task: Event-Customized Image Generation. It focuses on the customization of complex events with various target entities. Meanwhile, we proposed the first training-free event-customization framework FreeEvent. To facilitate this new task, we also collected two evaluation benchmarks from existing datasets and the internet, dubbed SWiG-Event and Real-Event, respectively. We validate the effectiveness of FreeEvent with extensive comparative and ablative experiments. Moving forward, we are going to 1) extend the event customization into other modalities, e.g., video generation; 2) explore advanced techniques for the finer combination of different customization works, e.g., subject, event, and style customizations.
Limitations. The main limitation of FreeEvent lies in the complexity of events and the number of entities. The customization effect may be compromised when there are too many entities in an image, especially if they are too small. As the first work in this direction, we hope our method can unveil new possibilities for more complex customization and the generation of a greater number of richer, more diverse entities. Additionally, since our model is built on pretrained Stable Diffusion (SD) models, our performance depends on the generative capabilities of SD. This can lead to suboptimal results for entities that the current SD struggles with, such as human faces and hands.
Potential Negative Societal Impacts. Since FreeEvent can seamlessly integrate with subject customization methods to generate target entities based on user-specified subjects, this capability also raises the same concerns about the potential misuse of pretrained SD models for malicious applications (e.g., Deepfakes) involving real human figures. To address this, it is essential to implement robust safeguards and ethical guidelines, similar to the security measures and NSFW content detection mechanisms already present in existing diffusion models.
@@ 2. $\bullet$
    In Sec. B, we show more details of the SWiG-Event benchmark and the process of retrieval-based quantitative comparisons.
@@ 3. $\bullet$
    In Sec. C, we show more qualitative comparison results of event customization on the Real-Event. .
@@
Appendix B Details of SWiG-Event and retrieval-based comparisons.
As shown in Figure 7(a), each SWiG-Event sample consists of a reference image with labeled bounding boxes and masks for each reference entity, the nouns of each reference entity, and the event class. As shown in Figure 7(b), we constructed the target prompt as a list of reference entity nouns. The ControlNet takes the semantic map merged from the masks as the layout condition, and BoxDiff takes the bounding boxes with labeled entity nouns as the layout condition. We then retrieved the target image for its corresponding reference image across all the 100 reference images that have the same reference event class.
Figure 7: (a) The SWiG-Event sample. (b) The process of the retrieval-based comparisons.
Appendix C More Qualitative Comparision Results.
We show more comparisons on Real-Event in Figure 8, Figure 9 and Figure 10.

### v2

@@ 4.1 Experimental Setup
Evaluation Benchmarks. In order to provide sufficient and suitable conditions for both quantitative and qualitative comparisons on this new task, we collect two new benchmarks111 Due to limited space, more details/results are in the Appendix. . 1) For quantitative evaluation, we present SWiG-Event, a benchmark derived from SWiG (pratt2020grounded) dataset, which comprises 5,000 samples with various events and entities, i.e., 50 kinds of different actions, poses, and interactions, where each kind of event has 100 reference images, and each reference image contains 1 to 4 entities with labeled bounding boxes and nouns. 2) For qualitative evaluation, we present Real-Event, which comprises 30 high-quality reference images from HICO-DET (chao2015hico) and the internet with a wide range of events and entities (e.g., animal, human, object, and their combinations). We further employ Grounded-SAM (kirillov2023segment; ren2024grounded) to extract the mask of each entity.
Baselines. We compared several kinds of SOTA baselines. For conditioned text-to-image generation baselines, we compared with training-based method ControlNet (zhang2023adding), MIGC (zhou2024migc), and training-free method BoxDiff (xie2023BoxDiff). For localized editing baselines, we compared with training-free methods PnP (tumanyan2023plug) and MAG-Edit (mao2024mag). For customization baselines, we compared with training-based methods Dreambooth (ruiz2023dreambooth) and ReVersion (huang2024reversion).
Implementation Details. We use Stable Diffusion v2-1-base as base model for all methods. Images are generated with a resolution of 512x512 on a NVIDIA A100 GPU1.
@@ 4.2 Quantitative Comparisons
We compare FreeEvent with state-of-the-art conditional text-to-image generation baselines ControlNet (zhang2023adding), MIGC (zhou2024migc), and BoxDiff (xie2023BoxDiff) on the SWiG-Event.
Setting. Each reference image in SWiG-Event contains reference entities together with labeled event class, bounding boxes, nouns, and their corresponding masks. Specifically, we construct the target prompt as a list of reference entity nouns, i.e., we ask all the methods to reproduce the event of the reference image with the same reference event and same reference entities. Additionally, ControlNet takes the semantic map merged from the masks as the layout condition. MIGC and BoxDiff take the bounding boxes with labeled entity nouns as the layout condition1.
Evaluation. Our evaluations follow the principle of whether generated images are aligned or similar with their reference images, and we apply multiple metrics to evaluate the customization quality of 5,000 target images from different perspectives. 1) Global image similarity: image retrieval performance and similarity scores. We retrieved each target image for its corresponding reference image based on the CLIP score across all the 100 reference images that have the same reference event class. Specifically, we extracted the image feature of each image through a pre-trained CLIP (radford2021learning) visual encoder and calculated the cosine similarities for image retrieval. Besides, we also used the CLIP-I, DreamSim (fu2023dreamsim), and DINO (oquabdinov2) scores to evaluate the image alignment of generated images with their reference images. 2) Event similarity: verb detection performance. We utilized the verb detection model GSRTR (cho2021gsrtr) which was trained on the SWIG dataset to detect the verb class of each generated image, and then calculated the detection accuracy based on the annotations of the reference images (i.e., whether the generated images and their reference images have the same verb class). 3) Entity similarity: CLIP-T (radford2021learning) score. We use the CLIP-T score to evaluate the text alignment of the generated images with text prompts. 4) Standard image generation metric. For a more comprehensive comparison, we used the FID (heusel2017gans) score to evaluate the overall quality of generated images.
@@
Results. As shown in Table 1, we can observe: 1) FreeEvent has better retrieval performance than both ControlNet, MIGC, and BoxDiff. This demonstrates that the target images generated by FreeEvent better preserve the overall characteristics of the reference event and entity. 2) FreeEvent also achieves the best verb detection performance, which indicates our method can better preserve the interaction semantics of the generated images. 3) FreeEvent further achieves superior performance over baselines across all similarity scores and standard image generation metrics, indicating our method can generate images with better qualities and alignment with both the reference images and texts. These results all demonstrate the effectiveness of FreeEvent for event customization.
@@ 4.3 Qualitative Comparisons
We compare FreeEvent with a wide range of SOTA baselines on the Real-Event1, including conditioned text-to-image generation method ControlNet (zhang2023adding) and BoxDiff (xie2023BoxDiff), localized image editing method PnP (tumanyan2023plug) and MAG-Edit (mao2024mag), image customization methods Dreambooth (ruiz2023dreambooth) and ReVersion (huang2024reversion).
@@
Figure 5: Ablations of the proposed paths and the target prompt. The “guidance” and “regulation” denote the cross-attention guidance and cross-attention regulation in entity switching path, respectively. The “injection” denotes the event transferring path.
Results. As shown in Figure 4, we can observe: 1) Conditional text-to-image generation models ControlNet and BoxDiff can only maintain the rough layout of each entity and struggle to capture the detailed action, pose, or interaction between different entities. And they both failed to match the generated entity with the desired target prompt. 2) For localized image editing methods PnP and MAG-Edit, while they can capture the reference event, they both struggle to accurately generate the target entities, and suffer from severe appearance leakage between each target entity (e.g., orange and strawberry in row three, tiger and lion in row five), and sometimes even failed to edit and output the original content. 3) The subject-customization model Dreambooth and the relation-customization model ReVersion both failed to generate satisfying results. As discussed before, these training-based methods require multiple reference images and are unable to learn the specific event when facing only one reference image. 4) Obviously, our FreeEvent successfully achieves the customization of various complex events with novel combinations of target entities. Meanwhile, the ControlNet and the localized image editing models tend to generate the target entities strictly matching the mask of their corresponding reference entities (e.g., bird in row two), which appears very incongruous. On the contrary, the entities generated by FreeEvent not only match the layout of the reference entity but also keep it harmonious. After all, while we use the reference entity mask to guide the generation of each target entity, the cross-attention guidance focuses on directing the overall layout of each target entity and doesn’t restrict detailed appearance, allowing for more diverse generation of target entities1.
@@ 4.4 Ablations
Influence of Different Target Prompts. Noteably, in our paper, the target prompt only contains the nouns of the target entities, we then run the ablations to analyze the influence of different descriptions (i.e., verb, background, style) in the target prompt for event customization.
@@
Combination of Event and Subject Customization. We further validate the ability of our framework to combine with subject customization methods to generate target entities with user-specified subjects, i.e., represented by identifier tokens. We took the Break-A-Scene model (avrahami2023break) to learn identifier tokens for subjects and replaced the Stable Diffusion models in Figure 2 with the fine-tuned one.
Results. As shown in Figure 6, FreeEvent can effectively generate various given subjects in specific events. Specifically, FreeEvent enables the flexible generation of a wide range of subject concepts (e.g., humans, regular objects, and backgrounds) and their combinations. These results demonstrated the great potential of our framework for Event-Subject customization.
Model Ours ControlNet BoxDiff PnP MAGEdit DreamBooth ReVersion
HJ 50 25 7 26 23 8 3
Table 2: Results of the user studies on the Real-Event.
@@ 4.5 User Study
Setting. We conducted user studies on Real-Event to further evaluate the effectiveness of FreeEvent. Specifically, we invited 10 experts and gave them a reference image, a target prompt, and seven target images generated by different models. They are asked to select at least one and up to three images that they believe demonstrate the best results in event customization, taking into account the generation effects of the events and entities, as well as the overall coherence of the images. We prepared 50 trials and asked the experts to give their judgments. The target image which got more than six votes is regarded as human judgment.
@@ 5 Conclusion
In this paper, we proposed a new task: Event-Customized Image Generation. It focuses on the customization of complex events with various target entities. Meanwhile, we proposed the first training-free event-customization framework FreeEvent. To facilitate this new task, we also collected two evaluation benchmarks from existing datasets and the internet, dubbed SWiG-Event and Real-Event, respectively. We validate the effectiveness of FreeEvent with extensive comparative and ablative experiments. Moving forward, we are going to 1) extend the event customization into other modalities, e.g., video generation; 2) explore advanced techniques for the finer combination of different customization works, e.g., subject, event, and style customizations.
Acknowledgements
This work was supported by the National Key Research & Development Project of China (2024YFB3312900), Key R&D Program of Zhejiang (2025C01128), an Fundamental Research Funds for the Central Universities. Long Chen was supported by the Hong Kong SAR RGC Early Career Scheme (26208924), the National Natural Science Foundation of China Young Scholar Fund (62402408), Huawei Gift Fund, and the HKUST Sports Science and Technology Research Grant (SSTRG24EG04).
Impact Statement
Since FreeEvent can seamlessly integrate with subject customization methods to generate target entities based on user-specified subjects, this capability also raises the same concerns about the potential misuse of pretrained SD models for malicious applications (e.g., Deepfakes) involving real human figures. To address this, it is essential to implement robust safeguards and ethical guidelines, similar to the security measures and NSFW content detection mechanisms already present in existing diffusion models.
@@ Acknowledgements
Figure 7: (a) The SWiG-Event sample. (b) The process of quantitative evaluation and image retrieval.
@@ 2. $\bullet$
    In Sec. B, we show more details of the SWiG-Event benchmark and the process of quantitative evaluation and image retrieval.
@@ 3. $\bullet$
    In Sec. C, we show quantitative results for path effectiveness.
  4. $\bullet$
    In Sec. D, we show the results for attribute generation during event customization.
  5. $\bullet$
    In Sec. E, we show more results for event-subject customization comparisions.
  6. $\bullet$
    In Sec. F, we provide more discussion about the hyperparameter setting.
  7. $\bullet$
    In Sec. G, we provide the discussion of our work’s limitations.
  8. $\bullet$
    In Sec. H, we show more qualitative comparison results of event customization on the Real-Event.
@@ 8. $\bullet$
The denoising process was set with 50 steps. For entity switching path, for all blocks and layers containing the cross-attention module, we apply the cross-attention guidance during the first 10 steps. And apply the cross-attention regulation during the whole 50 steps. For event transferring path, we perform spatial feature injection for block and layer at $\left\{\mathrm{decoder\,block\,1:}[\mathrm{layer}\,1\right]\}$ during the whole 50 steps. And perform self-attention injection for blocks and layers at $\left\{\mathrm{decoder\,block\,1:}[\mathrm{layer}\,1,2],\mathrm{decoder\,block\,2:}[\mathrm{layer}\,0,1,2],\mathrm{decoder\,block\,3:}[\mathrm{layer}\,0,1,2]\right\}$ during the first 25 steps. We set the classifier-free guidance scale to 15.0.
Appendix B Details of SWiG-Event and process of image retrieval.
As shown in Figure 7(a), each SWiG-Event sample consists of a reference image with labeled bounding boxes and masks for each reference entity, the nouns of each reference entity, and the event class. As shown in Figure 7(b), we constructed the target prompt as a list of reference entity nouns. The ControlNet takes the semantic map merged from the masks as the layout condition, and BoxDiff takes the bounding boxes with labeled entity nouns as the layout condition.
To compare the image retrieval performance, we retrieved the target image for its corresponding reference image across all the 100 reference images that have the same reference event class.
@@
FreeEvent 0.5771 0.2865 0.3877 0.3445
FreeEvent w/o guidance 0.5140 0.1058 0.2079 0.2979
FreeEvent w/o regulation 0.5459 0.1297 0.3011 0.3205
FreeEvent w/o injection 0.4945 0.0825 0.2694 0.3311
Table 3: Quantitative results for ablation study of two paths on Real-Event.
Figure 8: The results of attribute generation during event customization.
Appendix C Quantitative results for Path Effectiveness.
We ran ablations of two paths on Real-Event. We evaluated the CLIP-I, DreamSim, and DINO scores for image similarity between reference images and the generated images. We also evaluated the CLIP-T score for image-text similarity between text prompts and the generated images. Table 3 demonstrates the effectiveness of two paths.
Appendix D Attribute Generation Results.
In this paper, we didn’t explicitly model the attributes during generation. However, as the results are shown in Figure 8, since we can generate extra content for background and style by giving corresponding text descriptions, we thus tried to model the attributes by giving extra adjectives to the target prompt as an easy and natural exploration. Meanwhile, to ensure the accurate generation of the attributes, we applied the cross-attention guidance and regulation on each attribute using the mask of the entity they describe. As the results shown in Figure 8, our method successfully addresses the attributes of the corresponding entity (e.g., colors, materials, and ages). After all, while the attribute part is not the primary focus of our work, our approach shows potential and effectiveness in addressing it, and we would be happy to conduct further research in our future work.
Figure 9: Comparisions of Event-Subject Customization. Different colors and numbers show the associations between reference entities and their corresponding target prompts.
Figure 10: Comparisions of Event Customization with different hyperparameters settings. Different colors and numbers show the associations between reference entities and their corresponding target prompts.
Appendix E More Event-Subject Customization Comparisons.
In this section, we provide more event-subject customization comparisons with exisitng subject swapping and multi-subject customization methods, including Anydoor (chen2024anydoor), PhotoSwap (gu2023photoswap) and MS-Diffuison (wang2024ms). As shown in Figure 9, FreeEvent outperforms all other methods.
We need to clarify that the primary focus of this paper is event customization, while event-subject combined customization is only a potential capability of FreeEvent, rather than a key aspect we intend to emphasize or compare with existing methods. Moreover, FreeEvent serves as a plug-and-play framework for event-subject combined customization, making it unsuitable for direct compare with subject customization methods as their settings and applicable scenarios differ. We will provide further discussions and results on event-subject customization in future work.
Appendix F Discussion about hyperparameter Setting.
As the first work on event customization, our goal is to enable FreeEvent to perform high-quality event customization across diverse reference images using a unified set of hyperparameters (see Appendix Sec A). This default setting ensures faithful event transferring while allowing flexibility in target entity generation.
However, when there is a large shape discrepancy between the target entity and the reference entity, the layout information from the reference entity may undesirably affect the appearance of the target entity. This reflects a fundamental trade-off between event transferring and entity switching: prioritizing accurate event customization based on the reference image may lead to some compromise in the generation of the target entity. As the example shown in Figure 10, when the shape differences are significant (horse vs. dinosaur), the default setup may result in suboptimal generation (Target Image 1).
A straightforward solution is to adjust the parameters of the event transferring and entity switching paths. Specifically, enhancing the entity switching to emphasize the generation of the dinosaur, while slightly reducing the strength of event transferring to mitigate layout constraints from the reference entity.
As the shown in Figure 10, we provide an updated result (Target Image 2) by increasing the number of cross-attention guidance steps (from 10 to 15) and reducing the number of injection steps to 60% of the original. This rebalancing enables a more suitable trade-off for this case, resulting in more prominent dinosaur features (e.g., shorter front claws and upright back legs) while still preserving the core event structure.
This case study also demonstrates a practical and accessible way for users to adjust the trade-off between event transferring and entity switching according to their own customization needs. Looking ahead, we plan to explore more flexible and general solutions, such as adaptive parameter scheduling during generation or more explicit entity switching mechanisms, to further improve the controllability and diversity of target entities while maintaining event fidelity.
Appendix G Limitation.
The main limitation of FreeEvent lies in the complexity of events and the number of entities. The customization effect may be compromised when there are too many entities in an image, especially if they are too small. As the first work in this direction, we hope our method can unveil new possibilities for more complex customization and the generation of a greater number of richer, more diverse entities. Additionally, since our model is built on pretrained Stable Diffusion (SD) models, our performance depends on the generative capabilities of SD. This can lead to suboptimal results for entities that the current SD struggles with, such as human faces and hands.
Appendix H More Qualitative Comparision Results.
We show more comparisons on Real-Event in Figure 11, Figure 12, Figure 13, Figure 14 and Figure 15. Specifically, we list them by the order of entity numbers. And we use different combinations of target entities for the same reference image to generate diverse target images.
(The figures are in the next pages.)